OpenStudy (anonymous):

Select the counterexample that makes this conjecture false: For any real number x, x2 ≥ x. A.x = ½ B.x = 0 C.x = |3| D.x = -2

OpenStudy (zzr0ck3r):

is (1/2)^2 = (1/4) >= 1/2?

OpenStudy (anonymous):

YES

OpenStudy (zzr0ck3r):

1/4 is greater than or equal to 1/2?

OpenStudy (anonymous):

1/4 is greater than 1/2

OpenStudy (zzr0ck3r):

think about a dollar is 1/4 of a dollar (25 cents) greater than 1/2 of a dollar(50 cents)

OpenStudy (anonymous):

no 1/2 is greater than 1/4

OpenStudy (zzr0ck3r):

:) this one is your answer

OpenStudy (zzr0ck3r):

note: x = 0 does not work because of the equal sign.

OpenStudy (anonymous):

.thankyou

OpenStudy (zzr0ck3r):

np